# Languages Transferred Within the Encoder: On Representation Transfer in Zero-Shot Multilingual Translation

### v1

@@ 5.1 Setup
5.2 Results
Table 1 shows the experimental results of training from scratch. Our proposed methods always perform best in comparison with strong baselines, or are close to the best method. In supervised translations of three benchmarks of Europarl-15/TED-19/OPUS-100, LoLE shows divergent results of 0.13/-0.04/-0.22 on enx and 0.11/0.05/-0.02 on xen. Similarly, LCLR shows diverse results of -0.05/-0.01/-0.03 and 0.04/-0.01/0.07, respectively. Then, Both achieves the results of 0.18/-0.02/0.03 on enx and 0.12/0.12/-0.03 on xen. The results suggest that our model is comparable to Vanilla on supervised translations. Therefore, given that the improvement of zero-shot translation potentially reduces the quality of supervised translations Gu et al. (2019); Zhang et al. (2020a); Liu et al. (2021), our methods are stronger in avoiding the degeneration of supervised language pairs compared with other methods. In zero-shot translations, BOTH outperforms Vanilla 1.55/1.33/2.93 for BLEU and 0.58/1.02/3.12 for BERTScore. In an assessment of the qualities individually, LoLE achieves gains of 1.53/1.22/2.85 for BLEU and 0.54/0.91/3.34 for BERTScore, and LCLR improves 1.06/0.14/0.09 scores for BLEU and 0.39/0.13/0.06 scores for BERTScore. The results indicate that our proposed methods are orthogonal. Moreover, the significant testing shows that our methods substantially improve the zero-shot translation capacity in training from scratch.
@@ 6.2 Analysis of Improved Representation
To further verify our hypothesis, we measure representation transfer in the model incorporating our proposed methods. As shown in Figures 8(a) and 8(b), both scenarios exhibit improvements on (i). Meanwhile, Figures 8(c) and 8(d) indicate that the entanglement of representations among languages at the encoder is resolved. The evidence suggests that LoLE effectively enhances representation transfer in the encoder. Additionally, (ii) and (iii) in Figures 8(a) and 8(b) also achieve higher scores at lower layers of the encoder, which suggests that LoLE indeed makes lower layers of the encoder focus on surface-level information. By contrast, as shown in Figures 8(e) and 8(f), the more stable trend of (iv) in both scenarios suggests that LCLR can improve the capacity of lower layers of the decoder to distinguish languages to improve zero-shot translations. In addition, Appendix G provides the representational analysis for fine-tuning models, which proves that target language features are consistently beneficial in the encoder.
@@ 7 Related Works
Although representational analysis can guide improving the zero-shot translation capacity in MNMT, prior analyses have shown seemingly contradictory conclusions. Kudugunta et al. (2019) found that the encoder clusters representations of different translations based on target language features, suggesting that the encoder is language-specific. However, this conclusion is indirect because comparing real translation pairs would accumulate errors, even though Gu et al. (2019); Liu et al. (2021); Qu and Watanabe (2022) can support the validity. Our work extends this conclusion by using identity pairs to show that representations of source languages are transferred to the target subspace. Notably, Zhang et al. (2021); Pires et al. (2023) reached similar conclusions by incorporating learnable weights in additional language-specific modules. Another viewpoint is that an ideal encoder is considered to learn only language-agnostic representations Pham et al. (2019); Zhu et al. (2020); Pan et al. (2021); Gu and Feng (2022); Gao et al. (2023). Our representational analysis demonstrates that the general cross-lingual state does not exist in the encoder and the alignment occurs within the representational subspace of the target language. Additionally, our analysis supports that delimiting the language-specific representation boundaries at the decoder side is also important Yang et al. (2021); Chen et al. (2023).
@@ 8 Conclusion
We systematically investigated the representational issue of zero-shot translation deficiency in multilingual neural machine translation models. Our analyses show that the encoder transfers translation representations from the source language to the target language, and aligns different languages at the target language subspace. We applied engineering practices to verify our findings by proposing two orthogonal methods, which substantially improve the zero-shot translation capacity. Thus, our findings are significant for guiding the improvement of the transferability of multilingual representations.
@@ 9 Limitations
This work has two limitations. First, the identity pair is a proxy of language-specific representations, which are based on bi-directional training, i.e., each non-central language appears in the encoder and decoder together. Therefore, we designed an additional study to investigate the impact by retraining a model by eliminating nlen and enit so that it and nl appear only in the encoder and decoder, respectively based on the analysis in Section 3.2. Then, we conducted a comparison by taking de as the middle language to perform the role of identity pairs in analysis. As shown in Figure 9, the target language features keep the same trend as shown in Section 3.2 to support our conclusion again, but the influence of source language features increases relatively.
The second limitation is our investigation is based on adding a language tag specified to the target language at the beginning of the source sentence for the encoder. Although this is the de facto MNMT training strategy Johnson et al. (2017); Aharoni et al. (2019); Arivazhagan et al. (2019a); Gu et al. (2019); Pham et al. (2019); Wu et al. (2021); Yang et al. (2021); Pan et al. (2021); Qu and Watanabe (2022); Chen et al. (2023); Gu and Feng (2022); Gao et al. (2023), the current pre-trained models Fan et al. (2020); Liu et al. (2020); Team et al. (2022) are not based on the strategy. Our methodologies are also effective in mBART50 Liu et al. (2020) and M2M Fan et al. (2020) because they are based on adding a source language tag for the encoder, where we can measure the effectiveness of improving the target language information. However, NLLB Team et al. (2022) is based on two target language tags at the encoder and the decoder, respectively, whose representation transfer is not clear from the conclusions of this work. Thus, our future work is to investigate the representation transfer of NLLB.
@@ Appendix A Sentence-level SVCCA Score
Based on the definition of Section 2.1, we denote hidden representations of a sentence by }, where , equals to the length or from either the encoder or decoder, and is the model dimension. Additionally, the practical length is when is fed into the encoder because the encoder receives the input concatenated by and . Then, we derive the sentence-level representation using average pooling , which mainly represents the language features and semantics of the source sentence rather than syntactic information because positional information is reduced.
@@ Appendix B Detailed Information of Datasets
This work involves three datasets, i.e., Europarl-15, TED-19, and OPUS-100, where Europarl-15 and TED-19 are used in preliminary experiments.
Europarl-15 is collected from MMCR4NLP, which has high-quality sentences and each sentence in a language is one-to-one corresponding to other languages, i.e., all language-specific sets have parallel semantics Koehn et al. (2005); Dabre and Kurohashi (2019), including 15 European languages from 4 language families. Specifically, Germanic includes en, de, nl, da, Romance includes es, pt, it, ro, Slavic includes sl, bg, pl, cs, and Uralic includes fi, et, hu. The training and validation sets cover 28 supervised language pairs where English is the central language used to bridge the non-central languages. The test set consists of all language pairs, including 182 zero-shot language pairs in addition to supervised language pairs. Each pair in the training set comprises 189,310 sentences. The validation and test sets for each pair contain 2,000 sentences, and we measure SVCCA scores in the test sets.
In contrast to the aligned corpus, TED-19 consists of 19 languages, including en, ar, he, ru, ko, it, ja, zh, es, nl, vi, tr, fr, pl, ro, fa, hr, cs, de, which belong to various language families without parallel semantics, from TED Talks Ye et al. (2018). Each pair contains 103,093 to 214,111 sentences in training. The training set comprises 6,551,456 sentences in total, and the validation and test sets for each pair contain 2,000 sentences. Because of the unparallel semantics of TED-19, we align ar, he, zh, hr, vi, ja to obtain 967 sentences for measuring SVCCA scores. In addition, the reason why the number of languages is 19 is that, first, TED Talks have 20 high-resource languages, which are supported in M2M Fan et al. (2020) and mBART50 Liu et al. (2020). However, the tokenization of th is problematic, resulting in deprecating th.
OPUS-100 consists of 95 languages, 188 pairs, and 109.2 million samples in total Zhang et al. (2020a); Yang et al. (2021), where 90 pairs comprise 1 million samples and 56 pairs have more than 0.1 million samples. Different from Yang et al. (2021), we do not use the provided valid set for zero-shot translation pairs in training to avoid biases when assessing the transferability of multilingual representations.
@@ Appendix C Detailed Settings of Models
We use the Transformer Vaswani et al. (2017) from Fairseq Ott et al. (2019) as the model. For the configuration of models trained on Europarl-15 and TED-19, we follow Kudugunta et al. (2019) to set 6 encoder and decoder layers. In addition, we investigate models in Appendix E with 8 and 10 encoder layers. We adopt a shared vocabulary trained by SentencePiece Kudo and Richardson (2018) with 50,000 tokens for both the encoder and decoder. The model consists of 4 attention heads, embedding size of 512, inner size of 1024, dropout rate of 0.2, maximum learning rate of 0.0005 with the inverse square root schedule and 4,000 warmup steps, and label smoothing rate of 0.1. We set the batch size to 8,000 tokens per GPU, apply Adam Kingma and Ba (2017) as the optimizer, and set temperature sampling with Arivazhagan et al. (2019b). We train the model with 60 epochs for Europarl-15 and 30 epochs for TED-15, and finally average the top 5 checkpoints using the loss on the validation set. Compared with the basic configuration, the models trained on OPUS-100 have 8 attention heads, embedding size of 512, inner size of 2048, dropout rate of 0.1, and shared vocabulary size of 64,000. We train the model of OPUS-100 for 400,000 update steps with a batch size of 8,000 tokens per GPU for OPUS-100 and directly use the best checkpoint selected using the loss on the validation set. Furthermore, models with Europarl-15 and TED-19 are trained on 8 NVIDIA V100 GPUs, and models with OPUS-100 are trained on 4 NVIDIA A6000 GPUs by setting –update-freq to 2 in Fairseq to simulate 8 GPUs.
Three pre-trained models are utilized in our experiments. The first is M2M-418M Fan et al. (2020), a model trained on standard multilingual translation tasks, that supports translation across 100 languages. It is based on Transformer architecture, configured with 12 encoder and decoder layers, embedding size of 1024, inner size of 4096, and vocabulary size of 128,112, which results in a total of 418 million parameters. The second model, M2M-1.2B Fan et al. (2020), extends the M2M-418M by enlarging the number of layers to 24 and the inner size to 8192, which culminates in 1.2 billion parameters. The last model is mBART50 Liu et al. (2020), which was trained on monolingual corpora across 50 languages following Lewis et al. (2019) and preliminarily fine-tuned for multilingual translation tasks. It shares the same parameter setup as M2M-418M with a vocabulary size of 250,053, which consists of 611 million parameters. We conduct experiments on TED-19 because all covered languages are supported by these models.
@@ Appendix D Detailed Introductions of Figure 1
Appendix E Investigation on Models with 8 and 10 Encoder Layers
Figure 12 shows the scenarios of 8 and 10 encoder layers to confirm our conclusion, where all the cases for (i) consistently reflect the same tendency. Together with the increased number of layers, the increased final scores of (i) indicate the stronger target language features of ① presented for representations. Meanwhile, we also observe that the scores of (iii) are close to or even exceed those of (ii) at some layers both in Figures 2 and 12. This proves that the feature of the source language is not the primary factor for representations in the encoder because representations are transferred to the subspaces of target languages. Thus, this further investigation supports our conclusion that language transferred within the encoder is general.
Appendix F Selecting Hyper-Parameters
@@
Appendix G Analysis of Improved Representation for Fine-tuning Pre-trained Models
@@
Given the positive correlation shown in Section 6.1, we compute SVCCA scores in the same way as done in Section 3.2 and show the results in Table 3. Unlike Section 3.2, we equally consider the encoder and decoder because the encoder is only related to the source language and does not transfer representations to the target language in the training strategy of M2M Fan et al. (2020) and mBART50 Liu et al. (2020). Additionally, the different training strategy is the primary reason that F.T. shows the same scores in (i) and (iii) and keeps 100.0 in (ii). Alternatively, although the scores of (i), which reflect target language features, decrease in our methods, the scores of (ii) and (iii) also decrease. As a result, the differences between the scores of (i), (ii), and (iii) increase, that is, the relative importance of target language features increases. This result proves our statements in Sections 3.3 and 6.1 again that target language features are consistently beneficial in the encoder. On the other hand, the decoder side shows the same tendency as the encoder side. This fits our motivation in Section 4.2 to further improve the discriminating ability of lower layers of the decoder, although the training strategy of M2M and mBART50 has already provided a high capacity in discrimination for the decoder.
Appendix H Token-level Alignments in Other Cases

### v2

@@ 6.2 Analysis of Improved Representation
We measure representation transfer in the model incorporating our proposed methods to verify our findings further. As shown in Figures 9(a) and 9(b), both scenarios exhibit improvements on (i). Meanwhile, Figures 9(c) and 9(d) indicate that the entanglement of representations among languages at the encoder is resolved. The evidence suggests that LoLE effectively enhances representation transfer in the encoder. Additionally, (ii) and (iii) in Figures 9(a) and 9(b) also achieve higher scores at lower layers of the encoder, which suggests that LoLE indeed makes lower layers of the encoder focus on surface-level information. By contrast, as shown in Figures 9(e) and 9(f), the more stable trend of (iv) in both scenarios suggests that LCLR can improve the capacity of lower layers of the decoder to distinguish languages to improve zero-shot translations. In addition, Appendix F provides the representational analysis for fine-tuning models, which proves that target language features are consistently beneficial in the encoder.
@@ 7 Related Works
Prior studies on analyzing multilingual representation in Section 2.2 led to several effective methods in MNMT. Some works focused on updating and constraining the encoder to improve multilingual representations, and the findings in discrepancy mentioned in Section 2.2 led to two distinct approaches. First, Pan et al. (2021); Gu and Feng (2022); Gao et al. (2023); Bu et al. (2024) suggested regularizing the encoder for aligning semantic information across different source languages by introducing additional training objectives. Similarly, Pham et al. (2019); Zhu et al. (2020) explicitly modified the output form of the encoder to transfer the representation of the source sentence toward a language-agnostic state. Second, Gu et al. (2019); Liu et al. (2021); Sun et al. (2024) introduced specialized modeling constraints to improve the encoder to transfer source sentence representations to the target language without adding extra parameters, and Zhang et al. (2021); Pires et al. (2023) enhanced the representation of target language information by simply adding language-specific modules. Additionally, Yang et al. (2021); Qu and Watanabe (2022); Bu et al. (2024) focused on improving the target language representation on the decoder side or adding modules specified to the target language to the decoder. Given that the above works can all be encompassed within our analyses, we argue that this work offers insights for future improvements in MNMT. Specifically, enhancing the encoder to transfer source language representations into the target language subspace and align semantic information within those subspaces is the key to improving MNMT.
In addition, a critical factor of this work is the introduction of the identity pair as an analytical tool. Specifically, while identity pairs have been heuristically used in prior works Tiedemann and Scherrer (2019); Thompson and Post (2020); Bu et al. (2024), as an assumed indicator of language-specific representation states, they have not been subject to systematic or quantitative analysis. In contrast, we explicitly define, validate, and utilize identity pairs to probe representational properties in a controlled and measurable way. This not only strengthens the empirical basis of our conclusions but also constitutes an important methodological contribution of this work.
@@ 8 Conclusion
We systematically investigated the representational issue of zero-shot translation deficiency in multilingual neural machine translation models. Our analyses show that the encoder transfers translation representations from the source language to the target language, and aligns semantics across different source languages at the target language subspace. We applied engineering practices to verify our findings by proposing two orthogonal methods, which substantially improve the zero-shot translation capacity. Thus, our findings are significant for guiding the improvement of the transferability of multilingual representations.
@@ 9 Limitations
This work has two limitations. First, the identity pair is a proxy of language representations based on bi-directional training, i.e., each non-central language appears in the encoder and decoder together. Therefore, we designed an additional study to investigate the impact by retraining a model by eliminating nlen and enit so that it and nl appear only in the encoder and decoder, respectively, based on the analysis in Section 3.2. Then, we conducted a comparison by taking de as the middle language to perform the role of identity pairs in analysis. As shown in Figure 10, the target language features keep the same trend as shown in Section 3.2 to support our conclusion again, but the influence of source language features increases relatively.
The second limitation is our investigation is based on adding a language tag specified to the target language at the beginning of the source sentence for the encoder. Although this is the de facto MNMT training strategy Johnson et al. (2017); Aharoni et al. (2019); Arivazhagan et al. (2019a); Gu et al. (2019); Pham et al. (2019); Wu et al. (2021); Yang et al. (2021); Pan et al. (2021); Qu and Watanabe (2022); Chen et al. (2023a); Gu and Feng (2022); Gao et al. (2023), the current open-source models Fan et al. (2020); Tang et al. (2020); Team et al. (2022) are based on another strategy, i.e., adding a source language tag at the encoder side and adding a target language tag at the decoder side. Although, in Section 5, we have shown our proposed methods also benefit models with this strategy, this effectiveness is proved by empirical experiments. Thus, our future work is to investigate the representation transfer of this strategy to guide further improvements.
10 Further Considerations
Ethical Consideration
All datasets used in this work are public data, which are proven harmless. Moreover, this work is foundational research and is not tied to particular applications. Thus, there is no ethical risk existed in this work.
Sustainability statement
As noted in Appendix C, the GPU used in training individual models is A6000, which has an estimated carbon dioxide emission of approximately 0.13 kg per hour.666Measured by https://mlco2.github.io/impact/. Specifically, models trained on Europarl-15 and TED-19 required approximately 48 GPU hours, while models trained on OPUS-100 necessitated around 192 GPU hours.
@@ Appendix A Sentence-level SVCCA Score
Based on the definition of Section 2.1, we denote hidden representations of a sentence by }, where , equals to the length or from either the encoder or decoder, and is the model dimension. Additionally, the practical length is when is fed into the encoder because the encoder receives the input concatenated by and 777 plays the role of translation instruction instead of a token belonging to the target language with semantics, thus, this concatenation would not influence the measurement by mixing target language information into the sentence representation within the encoder.. Then, we derive the sentence-level representation using average pooling , which mainly represents the language features and semantics of the source sentence rather than syntactic information because positional information is reduced.
@@ Appendix B Detailed Information of Datasets
This work involves three datasets, i.e., Europarl-15, TED-19, and OPUS-100, where Europarl-15 and TED-19 are used in preliminary experiments. The training sets of those three datasets have different sizes, but the validation and test sets of a pair generally contain 2,000 translation instances. In preliminary experiments, we measure SVCCA scores in the test sets because those instances are unseen in the training.
Europarl-15 is collected from MMCR4NLP, which has high-quality translation instances and each instance in a language is one-to-one corresponding to other languages, i.e., all language-specific sets have parallel semantics Koehn et al. (2005); Dabre and Kurohashi (2019), including 15 European languages from 4 language families. Specifically, Germanic includes en, de, nl, da, Romance includes es, pt, it, ro, Slavic includes sl, bg, pl, cs, and Uralic includes fi, et, hu. The training and validation sets cover 28 supervised translation pairs where English is the central language used to bridge the non-central languages. The test set consists of all language pairs, including 182 zero-shot translation pairs in addition to supervised translation pairs. Finally, each pair in the training set comprises 189,310 instances.
In contrast to Europarl-15, which is the semantically aligned dataset, TED-19 consists of 19 languages, including en, ar, he, ru, ko, it, ja, zh, es, nl, vi, tr, fr, pl, ro, fa, hr, cs, de, which belong to various language families without parallel semantics, from TED Talks Ye et al. (2018). Each translation pair contains 103,093 to 214,111 instances in training, and the training set comprises 6,551,456 instances in total. Because of the unparallel semantics of TED-19, we align ar, he, zh, hr, vi, ja to obtain 967 translation instances for measuring SVCCA scores. In addition, the reason why the number of languages is 19 is that, first, TED Talks have 20 high-resource languages, which are supported in M2M Fan et al. (2020) and mBART50 Tang et al. (2020). However, the tokenization of th is problematic, resulting in deprecating th.
OPUS-100 consists of 95 languages, 188 pairs, and 109.2 million instances in total Zhang et al. (2020a); Yang et al. (2021), where 90 pairs comprise 1 million instances and 56 pairs have more than 0.1 million instances. Different from Yang et al. (2021), we do not include the zero-shot translation pairs in the validation set to avoid biases when assessing the transferability of multilingual representations.
@@ Appendix C Detailed Settings of Models
We implement the Transformer Vaswani et al. (2017) as the backbone model via Fairseq Ott et al. (2019). For the configuration of models trained on Europarl-15 and TED-19, we follow Kudugunta et al. (2019) to set 6 encoder and decoder layers. Based on the ablation study conducted in the validation set in Europarl-15 and TED-19 shown in Appendix E, we apply LoLE in the fifth encoder layer, set to 128, and set and of LCLR to 64 and 30, respectively. When we solely apply LCLR, we set the position to the bottom decoder layer based on the findings in Section 3.4. When we integrate both LoLE and LCLR into a model, we relocate LCLR to the second-bottom decoder layer because of the improved language features of the encoder representations. We adopt a shared vocabulary trained by SentencePiece Kudo and Richardson (2018) with 50,000 tokens for both the encoder and decoder. The model consists of 4 attention heads, embedding size of 512, inner size of 1024, dropout rate of 0.2, maximum learning rate of 0.0005 with the inverse square root schedule and 4,000 warmup steps, and label smoothing rate of 0.1. We set the batch size to 8,000 tokens per GPU, apply Adam Kingma and Ba (2017) as the optimizer, and set temperature sampling with Arivazhagan et al. (2019b). We train the model with 60 epochs for Europarl-15 and 30 epochs for TED-15, and finally average the top 5 checkpoints using the loss on the validation set. Compared with the basic configuration, the models trained on OPUS-100 have 8 attention heads, embedding size of 512, inner size of 2048, dropout rate of 0.1, and shared vocabulary size of 64,000. We enlarge to 256 and to 128 for models trained on OPUS-100 and three pre-trained models because they involve more languages. We train the model of OPUS-100 for 400,000 update steps with a batch size of 8,000 tokens per GPU for OPUS-100 and directly use the best checkpoint selected using the loss on the validation set. Furthermore, models with Europarl-15 and TED-19 are trained on 8 NVIDIA V100 GPUs, and models with OPUS-100 are trained on 4 NVIDIA A6000 GPUs by setting –update-freq to 2 in Fairseq to simulate 8 GPUs.
Three open-source models are utilized in fine-tuning experiments. The first is M2M-418M Fan et al. (2020), trained on standard multilingual translation tasks and supporting translation across 100 languages. It is based on Transformer architecture, configured with 12 encoder and decoder layers, embedding size of 1024, inner size of 4096, and vocabulary size of 128,112, which results in a total of 418 million parameters. The second model, M2M-1.2B Fan et al. (2020), enlarges the number of layers to 24 and the inner size to 8192 on M2M-418M, and culminates in 1.2 billion parameters. The last model is mBART50 Tang et al. (2020), trained on monolingual corpora across 50 languages following Lewis et al. (2019); Liu et al. (2020) and preliminarily fine-tuned for MNMT. It shares the same parameter setup as M2M-418M with a vocabulary size of 250,053, which consists of 611 million parameters. We conduct experiments on TED-19 because all covered languages are supported by these models.
@@ Appendix D Detailed Introductions of Figure 1
Appendix E Selecting Hyper-Parameters
@@
Appendix F Analysis of Improved Representation for Fine-tuning Pre-trained Models
@@
Given the positive correlation shown in Section 6.1, we compute SVCCA scores in the same way as done in Section 3.2 and show the results in Table 3. Unlike Section 3.2, we equally consider the encoder and decoder because the encoder is only related to the source language and does not transfer representations to the target language in the training strategy of M2M Fan et al. (2020) and mBART50 Tang et al. (2020). Additionally, the different training strategy is the primary reason that F.T. shows the same scores in (i) and (iii) and keeps 100.0 in (ii). Alternatively, although the scores of (i), which reflect target language features, decrease in our methods, the scores of (ii) and (iii) also decrease. As a result, the differences between the scores of (i), (ii), and (iii) increase, that is, the relative importance of target language features increases. This result proves our statements in Sections 3.3 and 6.1 again that target language features are consistently beneficial in the encoder. On the other hand, the decoder side shows the same tendency as the encoder side. This fits our motivation in Section 4.2 to further improve the discriminating ability of lower layers of the decoder, although the training strategy of M2M and mBART50 has already provided a high capacity in discrimination for the decoder.
Appendix G Token-level Alignments in Other Cases
@@
Appendix H Evaluation Metrics Selection
In this work, we select two main automatic evaluation metrics and a secondary statistic measurement. The first one is SacreBLEU Post (2018) which is an implementation of BLEU Papineni et al. (2002). This is the most popular and common metric used in evaluating the alignment between inferences and references at the word level. In order to counter the insufficiency of SacreBLEU, we also select BERTScore Zhang et al. (2020b), which is a representational metric to evaluate the semantic similarity between inferences and references. Furthermore, to show whether the improvements brought by proposed methods are significant, we also conduct the statistical significance testing Koehn (2004) using paired bootstrap resampling with 1,000 iterations and 0.5 resampling ratios, consequently, the case of 0.05 means that the difference is significant.
Additionally, we follow prior works Yang et al. (2021); Chen et al. (2023a) to report the off-target ratio, which is measured by fasttext-langdetect888https://pypi.org/project/fasttext-langdetect. The off-target translation refers to a sentence translated to an incorrect target language rather than the target language we expected. However, the off-target ratio is not reliable, because the popular tools used in measuring off-target ratios are based on word level and lack support in low-resource languages. Furthermore, the score of SacreBLEU can directly show the problem of off-target, because the evaluation process of SacreBLEU tends to give a great penalty on an inference, which has a different writing script from the expected target language. Therefore, we only report it as a secondary metric in Appendix I.
Appendix I Off-Target Ratio of Results
Europarl-15 | TED-19 | OPUS-100 | ||||
Method | zero.() | off.() | zero.() | off.() | zero.() | off.() |
Vanilla | 24.65 | 1.34 | 11.98 | 4.08 | 5.04 | 70.41 |
DisPos | 25.89 | 0.84 | 12.80 | 3.82 | 5.58 | 61.65 |
TLP | 24.96 | 1.22 | 12.74 | 3.71 | 4.60 | 83.29 |
SemAli | 25.25 | 0.99 | 13.45 | 3.62 | 6.42 | 58.25 |
LoLE | 26.09 | 0.71 | 13.20 | 3.69 | 7.92 | 50.05 |
LCLR | 25.71 | 0.79 | 12.12 | 3.86 | 5.11 | 68.53 |
BOTH | 26.20 | 0.74 | 13.31 | 3.69 | 7.97 | 55.06 |
Model | Metric | Pre. | F.T. | LOLE | LCLR | BOTH |
M2M-418M | zero.() | 14.51 | 17.46 | 17.52 | 17.65 | 17.68 |
off.() | 3.66 | 3.34 | 3.32 | 3.24 | 3.33 | |
M2M-418M | zero.() | 15.95 | 18.48 | 18.67 | 18.64 | 18.69 |
off.() | 3.50 | 3.15 | 3.16 | 3.16 | 3.14 | |
mBART50 | zero.() | 6.92 | 5.58 | 7.28 | 9.69 | 9.55 |
off.() | 43.56 | 65.26 | 40.76 | 38.24 | 35.28 |
Tables 4 and 5 show the measurement of the off-target ratio, which are the supplement of Tables 1 and 2. We can observe that the off-target ratio is always inversely proportional to BLEU scores, aligning with our discussion in Appendix H. Additionally, there are two points worth noting: (1) In Table 4, the off-target ratio in OPUS-100 is generally higher. This is not an outlier because resulting in a strong zero-shot translation capability in OPUS-100 is particularly challenging due to the large number of languages involved and the limited corpus for individual languages Zhang et al. (2020a); Yang et al. (2021). (2) In Table 5, the off-target ratio counted from mBART50 is higher than other cases. This abnormal value has been discussed in Section 5.2, that is, the zero-shot ability of mBART50 is weaker than M2M models, and then, the fine-tuning dramatically changes the behavior of the model.